OpenStudy (anonymous):

write an equation in point slope form of the line that passes through (6,3) and (-2,1)

OpenStudy (anonymous):

write an equation in point slope form of the line that passes through (6,3) and (-2,1)

OpenStudy (anonymous):

write an equation in point slope form of the line that passes through (6,3) and (-2,1)

OpenStudy (anonymous):

write an equation in point slope form of the line that passes through (6,3) and (-2,1)

OpenStudy (anonymous):

pt slope form is: y-y1=m(x-x1) so 6,3 =x1, y1 and (-2,1) =x2, y2 you can find the slope by doing the slop equation: where slpe =y2-y1/x2-x1, that will give your the slop (m) you need for for your point slope

OpenStudy (anonymous):

sorry, i just realized that this a 2 point form. so you will need the equation: y-y1=y2-y1/x2-x1(x-x1)

OpenStudy (anonymous):

it should look like this: y-3=(1-3/-2-6)(x-6)